# PHANGS-JWST: Data Processing Pipeline and First Full Public Data Release

### v1

@@ 3.6.1 Single-Tile Destriping
We show an example of this single-tile destriping in Figure 4. This step effectively removes the $1/f$ noise on small scales whilst retaining the large-scale structure present in the images. However, large-scale ‘ripples’ remain, which we mitigate with our second destriping technique (Sect. 3.6.2).
@@
Figure 5: Applying the single-tile destriping to Spitzer observations of M33. The left and middle panels show the original and destriped data respectively, and are locked at the same colorscale. The right panel shows the difference image, on the order of 0.001 MJy/sr, essentially below the noise level of the original image.
@@ 3.6.2 Multi-Tile Destriping
In detail, for each exposure, we create a sigma-clipped median image of all overlapping exposures. For PHANGS-JWST, there are at least four overlapping exposures for all bands and up to 12 for F200W. When creating this image, we make sure that the overall flux level matches between tiles by performing this after our background matching step (Sect. 3.5). For each individual exposure, we then create a difference image by comparing that exposure to this median image. Because it is present in both the median and the image, real structure will not be present in this difference image but it will include any large-scale ripples. Using this difference image we then calculate a sigma-clipped median along rows, per-amplifier as before. We subtract these medians from our image to effectively remove any large-scale stripes still left in the data. An example of this procedure is shown in Figure 6. We also attempted to perform destriping using only this method, but found that the small-scale stripes remained; hence, both the single- and multi-tile destriping are necessary to remove the $1/f$ noise on all scales.
Particularly in the short NIRCam observations, there remains some very low-level striping in the data following this first pass using overlapping tiles. This reflects the fact that whilst the $1/f$ noise is essentially random, given its prevalence it can interfere constructively when creating an average image. We therefore carry out a second pass where we create a smoothed, stacked image, performing a large-scale (a tenth the size of the full array) median filter along the axis perpendicular to the direction of the stripes in the stacked image. This creates a new median image as in the previous step, but with any remaining large-scale noise filtered out. We then subtract the large-scale median filtered image from the individual tiles and then take a row-by-row median to construct a last noise estimate. We calculate the median across the full tile rather than per amplifier, as the median filter redistributes flux, potentially leading to a bias along rows completely full of emission. We show an example of this final step in Figure 7, and note that these last remaining stripes have significantly smaller amplitude than those subtracted in the previous destriping steps. Whilst we have not performed detailed tests on how much our destriping lowers the noise in our imaging, initial tests in blank fields appear to lower the RMS noise level by around 10 percent, consistent with other algorithms (e.g. Bagley et al., 2023). We leave a full exploration of optimal destriping methods and noise improvements to future work. Having carried out these destriping steps, the images appear stripe-free.
@@ 3.7 Level Three Processing
We perform outlier flagging with default parameters, although to speed this step up we keep all images in memory (in_memory = True; if this setting is False then this step takes on the order of hours, rather than minutes). F200W is the most memory-intensive band here, and can use up to 250GB of RAM. Following Level Three processing, we have final mosaics in all of our eight bands, which are aligned north-up.
@@ 3.8 Absolute Astrometric Alignment
For PHANGS-JWST, these are luminous red stellar sources, a mix of AGB stars and red supergiants, that are bright in our JWST bands. These stars are identified from HST imaging, and their astrometric calibration descends from those data (Lee et al., 2022). To select these sources, we use a combination of detection quality and color-magnitude cuts. We require ${\rm S/N_{F814W}}\geq 10$, ${\rm crowd_{F814W}}\leq 0.01$, ${\rm PhotQualFlag_{F814W}}=0$, ${\rm OBJTYPE}=1$, $-0.025<{\rm sharp_{F814W}}\leq 0.01$ and $-0.15<{\rm round_{F814W}}\leq 0.4$ to obtain bright, uncrowded, well-PSF-fitted point-like sources. For descriptions of these various parameters, we refer the reader to Thilker et al. (2022) and the DOLPHOT (Dolphin, 2000, 2016) documentation (http://americano.dolphinsim.com/dolphot/) Sharpness and roundness limits are slightly adjusted if necessary per galaxy. For the color-magnitude cuts, we require F814W in the range $3-5$ mag brighter than the foreground extinction corrected tip of the red giant branch (TRGB) and take only sources with $V$-$I$ color in the range $42\leq F555W-F814W<$ 4 mag.
@@
We perform this procedure separately for each NIRCam band. For MIRI, there are a lack of true point sources at many of the longer wavelength bands (particularly 21 µm). For this, we calculate a solution from the F1000W band (as the highest resolution MIRI band with many clear point sources), and apply this solution to all the MIRI bands as they have consistent internal astrometry. At this stage, we also back-propagate these astrometric solutions to the individual tiles, as these are used as the inputs to the stellar photometry package DOLPHOT (Dolphin, 2000, 2016), which can be used for cluster photometry via PSF fitting.
@@ 3.9 Anchoring the Background Level
[…]
@@
[…]
@@ 3.10 PSF Matching
FWHM [${}^{\prime\prime}$] FWHM [${}^{\prime\prime}$] FWHM [${}^{\prime\prime}$] FWHM [${}^{\prime\prime}$]
@@
Table 1: FWHM for very safe, safe, and aggressive convolution kernels for our JWST filters, calculated using the metrics of Aniano et al. (2011).
@@
Users interested in mitigating the impact of the JWST PSF may wish to convolve the data with a suitable kernel to obtain the smallest “safe” Gaussian PSF for a specific band, different from our default F2100W. We report in Table 1 the Gaussians PSFs corresponding to “very safe”, “safe”, and “aggressive” kernels according to Aniano et al. (2011). Suitable kernels for all these bands can be generated using our jwst_kernels code with the make_jwst_kernel_to_Gauss function.
@@ 4 Data Quality
Figure 9: Final mosaics for each band for NGC 628. NIRCam filters are shown in the top row, and MIRI on the bottom. Each image is on a linear stretch between the $2^{\rm nd}$ and $98^{\rm th}$ percentiles of the pixel distribution, to highlight the fainter structure in the images.
@@ 4.2 Noise Estimates Compared to Error Maps
Using the background regions from the previous Subsection, we compare the error measured at the native resolution to the pipeline-calculated error map. We take the standard deviation of the pixels within the background region in the science image for the four MIRI bands (we do not use NIRCam as due to the source density, we are unable to find any background regions within our images), and compare these to the mean of the pixels within the same region of the error map image. This is shown in Figure 11 – there is a good agreement here, with the error map value typically being higher by around 5 to 10 percent. We therefore deem these error maps suitable for providing estimates of the noise for observations where no background regions are present.
@@ 4.3 Quality Assurance and Outstanding Issues
Having created our final mosaics, we perform a round of internal quality assurance (QA). We have examined each individual image for issues such as poor astrometric alignment (which manifests as “blurring” between tiles), stepping between tiles which indicates poor level matching, and other defects in the data quality (e.g., remaining $1/f$ noise). We find that each mosaic is free of these issues, and the data quality is excellent.
We then performed a round of inter-band quality checks, producing three-color images (of F200W, F300M, and F1000W) using SAOImageDS9 (Joye & Mandel, 2003) to assess the NIRCam short, NIRCam long, and MIRI overall alignment. Here, we are essentially checking that the astrometry between these mosaics is good, as large misalignments will lead to clear blurring between the color channels. We find that the NIRCam-to-NIRCam agreement is excellent, as is the NIRCam-to-MIRI agreement in most cases, although there are some exceptions. The mosaics themselves generally look excellent, but there are some outstanding issues that remain unresolved in this data release. We expect that these may improve as characterization of the instrumental features of JWST improve, and substantial improvements will be followed by new version releases. In particular, our QA has noted the following issues:
@@ 4.3.1 PSF spikes and saturation in bright galaxy centers
A few bright sources lead to large PSF spikes that can cover a significant portion of the observations. These spikes primarily affect MIRI images and they mostly original from bright point sources, both star clusters and AGN, in galaxy centers. We have experimented with using WebbPSF to subtract these and so recover the affected extended emission. However, we find that WebbPSF models matched to the observed sources tend to oversubtract the spikes, particularly near the bright sources themselves. This could be due to the “brighter-fatter” effect (Argyriou et al., 2023) whereby charge can leak into neighbouring pixels, leading to an increase in the PSF size of 10–25 percent in bright sources. Additionally, limitations in the current theoretical model for the PSF are likely prevalent in the outskirts of the PSF wings. Whatever the root cause, we currently do not have a robust method for removing these PSF spikes. Lacking the ability to correct for these, many of our analyses identify and mask regions affected by PSF spikes before analyzing the data.
@@ 4.3.2 Astrometry issues
In Figures 12 and 13, we show the quality of the alignment using sources confidently matched by eye between HST, NIRCam, and MIRI observations. Whilst we can see that the NIRCam-to-HST alignment is generally very good – the median offset across all NIRCam bands with respect to HST is $\farcs$013 (around 0.5 NIRCam short or 0.2 NIRCam long pixels) with a median RMS scatter of $\farcs$007. However, the agreement is less good between MIRI and NIRCam – the median offset across all MIRI bands with respect to F300M is $\farcs$06 (around 0.5 MIRI pixels) with a median RMS scatter of $\farcs$04. This is mainly driven by the much lower density of point sources in the MIRI data, and the source extraction not being optimized for the JWST PSF shape. As well as this, in some cases, a point source at the MIRI resolution (particularly F2100W) can be an extended source at the higher NIRCam resolution (often a background galaxy), making this comparison more complicated. The shifts are typically less than a pixel but can be more severe in some cases; our worst case is NGC3351, where the shift between NIRCam and MIRI is around one MIRI pixel. In a few cases there are multiple peaks in the distribution of residual astrometric offsets, indicating some persistent unresolved issues with the tile-to-tile matching.
@@ 5 Existing mosaic comparison
In this Section, we show the level of changes in our data processing relative to both our early science data release and the mosaics obtainable through the MAST service. Overall, our latest version (DR1.0.1) is a significant improvement over both of these products, and so we suggest users to prefer our latest versions over these other available versions.
@@ 5.1 Comparison to early science release
Figure 14: Comparison between our latest reprocessing and our early science mosaic for NGC 628. The difference here is the early science version subtracted from our latest reprocessed version, after reprojecting to the same WCS. There are significant differences in the short NIRCam band, caused by both background level matching and our new destriping method. At long NIRCam wavelengths, the differences are relatively minor. For MIRI, the improved calibration shows up as a galaxy imprint (most clear at F770W). Differences in astrometry show neighboring red/blue shifts and the background matching algorithm shows clear steps between mosaic tiles.

### v3

@@ 3.6.1 Single-Tile Destriping
We show an example of this single-tile destriping in Figure 4. We find the amplitude of the noise removed in this step to be typically around the 0.05 DN s-1 level. This step effectively removes the $1/f$ noise on small scales whilst retaining the large-scale structure present in the images. However, large-scale ‘ripples’ remain, which we mitigate with our second destriping technique (Sect. 3.6.2).
@@
Figure 5: Applying the single-tile destriping to Spitzer observations of M33. The left and middle panels show the original and destriped data respectively, and are locked at the same colorscale. The right panel shows the difference image, on the order of 0.001 MJy/sr, essentially below the noise level of the original image. The white pixels in the difference image are due to pixels with no data; these are not present in the JWST noise models but are present here due to differences in the specifics of the data between JWST and Spitzer.
@@ 3.6.2 Multi-Tile Destriping
In detail, for each exposure, we create a sigma-clipped median image of all overlapping exposures. For PHANGS-JWST, there are at least four overlapping exposures for all bands and up to 12 for F200W. When creating this image, we make sure that the overall flux level matches between tiles by performing this after our background matching step (Sect. 3.5). For each individual exposure, we then create a difference image by comparing that exposure to this median image. Because it is present in both the median and the image, real structure will not be present in this difference image but it will include any large-scale ripples. Using this difference image we then calculate a sigma-clipped median along rows, per-amplifier as before. We subtract these medians from our image to effectively remove any large-scale stripes still left in the data. An example of this procedure is shown in Figure 6. Typically, the amplitude of the noise removed here is at around the 0.05 MJy sr-1 level, and is worst for the F200W band. We also attempted to perform destriping using only this method, but found that the small-scale stripes remained; hence, both the single- and multi-tile destriping are necessary to remove the $1/f$ noise on all scales.
Particularly in the short NIRCam observations, there remains some very low-level striping in the data following this first pass using overlapping tiles. This reflects the fact that whilst the $1/f$ noise is essentially random, given its prevalence it can interfere constructively when creating an average image. We therefore carry out a second pass where we create a smoothed, stacked image, performing a large-scale (a tenth the size of the full array) median filter along the axis perpendicular to the direction of the stripes in the stacked image. This creates a new median image as in the previous step, but with any remaining large-scale noise filtered out. We then subtract the large-scale median filtered image from the individual tiles and then take a row-by-row median to construct a last noise estimate. We calculate the median across the full tile rather than per amplifier, as the median filter redistributes flux, potentially leading to a bias along rows completely full of emission. We show an example of this final step in Figure 7, and note that these last remaining stripes have significantly smaller amplitude than those subtracted in the previous destriping steps, at around the 0.01 MJy sr-1 level. Whilst we have not performed detailed tests on how much our destriping lowers the noise in our imaging, initial tests in blank fields appear to lower the RMS noise level by around 10 percent, consistent with other algorithms (e.g. Bagley et al., 2023). We leave a full exploration of optimal destriping methods and noise improvements to future work. Having carried out these destriping steps, the images appear stripe-free.
@@ 3.7 Level Three Processing
We perform outlier flagging with default parameters, although to speed this step up we keep all images in memory (in_memory = True; if this setting is False then this step takes on the order of hours, rather than minutes). F200W is the most memory-intensive band here, and can use up to 250GB of RAM. Following Level Three processing, we have final mosaics in all of our eight bands, which are aligned north-up. These mosaics are produced at the native pixel scale of the data (i.e. $\farcs$031 for NIRCam short, $\farcs$063 for NIRCam long, $\farcs$11 for MIRI). We use the default pixfrac of 1.0, and note that as we perform another round of astrometric alignment following the mosaicking (Sect. 3.8), the images are not aligned to the same pixel grid. We expect that future versions of the data will feature products aligned to the same pixel grid.
@@ 3.8 Absolute Astrometric Alignment
For PHANGS-JWST, these are luminous red stellar sources, a mix of AGB stars and red supergiants, that are bright in our JWST bands. These stars are identified from DOLPHOT (Dolphin, 2000, 2016) catalogs produced from the HST imaging, and their astrometric calibration descends from those data (Lee et al., 2022). To select these sources, we use a combination of detection quality and color-magnitude cuts. To obtain bright, uncrowded, well-PSF-fitted point-like sources, we use the following DOLPHOT parameter criteria: ${\rm S/N_{F814W}}\geq 10$, ${\rm crowd_{F814W}}\leq 0.01$, ${\rm PhotQualFlag_{F814W}}=0$, ${\rm OBJTYPE}=1$, $-0.025<{\rm sharp_{F814W}}\leq 0.01$ and $-0.15<{\rm round_{F814W}}\leq 0.4$. The crowding parameter describes how much brighter a star would be (in magnitudes) if the flux from neighboring stars was not removed. By definition the value is zero for an isolated star. The object type is a classification for the morphology of the sources, where 1 corresponds to a well-detected star. The quality flag provides information on artifacts, such as bad or saturated pixels, and whether the extent of the source is not fully captured because it is at the edge of the chip. A flag of zero indicates no such issues. Sharpness is a morphological parameter that is computed relative to a point source: if a source is perfectly fit by the PSF model, then the value is zero. Positive values indicate sources are more centrally peaked than the PSF (e.g., a cosmic ray), while negative values indicate those that are more extended. Finally, roundness is a symmetry parameter, with a value of zero for perfectly circular sources. For further discussion, we refer the reader to Thilker et al. (2022) and the DOLPHOT (Dolphin, 2000, 2016) documentation (http://americano.dolphinsim.com/dolphot/) Sharpness and roundness limits are slightly adjusted if necessary per galaxy. For the color-magnitude cuts, we require F814W in the range $3-5$ mag brighter than the foreground extinction corrected tip of the red giant branch (TRGB) and take only sources with $V$-$I$ color in the range $42\leq F555W-F814W<$ 4 mag.
@@
We perform this procedure separately for each NIRCam band. For MIRI, there are a lack of true point sources at many of the longer wavelength bands (particularly 21 µm). For this, we calculate a solution from the F1000W band (as the highest resolution MIRI band with many clear point sources), and apply this solution to all the MIRI bands as they have consistent internal astrometry. At this stage, we also back-propagate these astrometric solutions to the individual tiles, as these are used as the inputs to the stellar photometry package DOLPHOT (Dolphin, 2000, 2016).
@@ 3.9 Anchoring the Background Level
After the tile background matching step above, our mosaic tiles have a consistent internal background (Sect. 3.5), with little or no jump in intensity from tile to tile within a mosaic. However, there is still uncertainty in the overall intensity of the background level. That is, at this stage the intensity of empty sky near the galaxy is not necessarily zero, and our empirical tests suggest background level variations by several times $\sim 0.1$ MJy sr-1 from galaxy to galaxy and band to band to be common, even up to ±$\sim\pm 1$ MJy sr-1 in some cases. Given the small field of view of many PHANGS-JWST mosaic compared to the angular extent of the galaxy, it is often not trivial to identify a region of true galaxy-free sky to establish the background level.
@@
The procedure appears to work empirically. The bands show tight, clear linear relations with one another, almost always similar to Fig. 8. The ratio images among bands in low intensity parts of galaxies, which tend to be sensitive to background-level discrepancies, also appear mostly free of background-related artifacts. Most directly, in a few cases we do have some regions that contain mostly empty sky (which we will also use to study the noise properties as a function of resolution in Sect. 4.1). After this anchoring procedure, the intensity histograms appear centered at intensities within a few times $\sim 0.01$ MJy sr-1 of zero.
We caution that the anchoring to a true, correct sky background will only be as good as the reference image. Our IRAC1 images come mostly from S4G (Sheth et al., 2010), our IRAC4 images from SINGS (Kennicutt et al., 2003) or LVL (Dale et al., 2009), and our WISE1 and WISE3 images from z0MGS (Leroy et al., 2019). In each case, we conducted an additional round of by-hand background fitting to the reference image, but we do note that this remains an area where additional work (as well as improvements to the algorithm) may further improve the images at the few times $0.01$ MJy sr-1 level.
@@ 3.10 PSF Matching
FWHM [′′] FWHM [′′] FWHM [′′] FWHM [′′]
@@
Table 2: FWHM for very safe, safe, and aggressive convolution kernels to a Gaussian PSF for our JWST filters, calculated using the metrics of Aniano et al. (2011).
@@
Users interested in mitigating the impact of the JWST PSF may wish to convolve the data with a suitable kernel to obtain the smallest “safe” Gaussian PSF for a specific band, different from our default F2100W. We report in Table 2 the Gaussians PSFs corresponding to “very safe”, “safe”, and “aggressive” kernels according to Aniano et al. (2011). Suitable kernels for all these bands can be generated using our jwst_kernels code with the make_jwst_kernel_to_Gauss function. For matched PSF work, we recommend experimenting with the kernels presented here; if resolution is particularly important, then opting for an “aggressive” kernel may be the optimal choice. However, these more aggressive convolutions are more prone to artifacts, especially around very bright sources. We therefore strongly suggest using visual inspection to ensure any convolved images are artifact free to the level the science requires.
@@ 4 Data Quality
Figure 9: Final mosaics for each band for NGC 628. NIRCam filters are shown in the top row, and MIRI on the bottom. Each image is on a linear stretch between the $2^{\rm nd}$ and $98^{\rm th}$ percentiles of the pixel distribution, to highlight the fainter structure in the images. The complete figure set (19 images) is available in the online journal.
@@ 4.2 Noise Estimates Compared to Error Maps
Using the background regions from the previous Subsection, we compare the error measured at the native resolution to the pipeline-calculated error map. We take the standard deviation of the pixels within the background region in the science image for the four MIRI bands (we do not use NIRCam as due to the source density, we are unable to find any background regions within our images, although given the good agreement between our calculated sensitivities and the exposure time calculator estimates, these error maps are likely also reasonable), and compare these to the mean of the pixels within the same region of the error map image. We note that this error map does not account for pixel-correlated noise. We have not included this effect at this point, but will revisit this at some point in the future. For an example of how this can be accounted for, we refer readers to Yang et al. (2023). This comparison is shown in Figure 11 – there is a good agreement here, with the error map value typically being higher by around 5 to 10 percent. We therefore deem these error maps suitable for providing estimates of the noise for observations where no background regions are present.
@@ 4.3 Quality Assurance and Outstanding Issues
Having created our final mosaics, we perform a round of internal quality assurance (QA). We have examined each individual image for issues such as poor astrometric alignment (which manifests as “blurring” between tiles), stepping between tiles which indicates poor level matching, and other defects in the data quality (e.g., remaining $1/f$ noise). We find that each mosaic is free of these issues.
We then performed a round of inter-band quality checks, producing three-color images (of F200W, F300M, and F1000W) using SAOImageDS9 (Joye & Mandel, 2003) to assess the NIRCam short, NIRCam long, and MIRI overall alignment. Here, we are essentially checking that the astrometry between these mosaics is good, as large misalignments will lead to clear blurring between the color channels. We find that the NIRCam-to-NIRCam agreement is very good (with any misalignment significantly less than a NIRCam pixel), as is the NIRCam-to-MIRI agreement in most cases (with any misalignment less than a MIRI pixel), although there are some exceptions. The mosaics themselves generally look excellent and are ready for a wide range of science cases, but there are some outstanding issues that remain unresolved in this data release. We expect that these may improve as characterization of the instrumental features of JWST improve, and substantial improvements will be followed by new version releases. In particular, our QA has noted the following issues:
@@ 4.3.1 PSF spikes and saturation in bright galaxy centers
A few bright sources lead to large PSF spikes that can cover a significant portion of the observations. These spikes primarily affect MIRI images and they mostly originate from bright point sources, both star clusters and AGN, in galaxy centers. We have experimented with using WebbPSF to subtract these and so recover the affected extended emission. However, we find that WebbPSF models matched to the observed sources tend to oversubtract the spikes, particularly near the bright sources themselves. This could be due to the “brighter-fatter” effect (Argyriou et al., 2023) whereby charge can leak into neighbouring pixels, leading to an increase in the PSF size of 10–25 percent in bright sources. Additionally, limitations in the current theoretical model for the PSF are likely prevalent in the outskirts of the PSF wings. Whatever the root cause, we currently do not have a robust method for removing these PSF spikes. Lacking the ability to correct for these, many of our analyses identify and mask regions affected by PSF spikes before analyzing the data.
@@ 4.3.2 Astrometry issues
In Figures 12 and 13, we show the quality of the alignment using sources confidently matched by eye between HST, NIRCam, and MIRI observations. Whilst we can see that the NIRCam-to-HST alignment is generally very good – the median offset across all NIRCam bands with respect to HST is $\farcs$012 (around 0.5 NIRCam short or 0.2 NIRCam long pixels) with a median RMS scatter of $\farcs$006. However, the agreement is less good between MIRI and NIRCam – the median offset across all MIRI bands with respect to F300M is $\farcs$05 (around 0.5 MIRI pixels) with a median RMS scatter of $\farcs$03. This is mainly driven by the much lower density of point sources in the MIRI data, and the source extraction not being optimized for the JWST PSF shape. As well as this, in some cases, a point source at the MIRI resolution (particularly F2100W) can be an extended source at the higher NIRCam resolution (often a background galaxy), making this comparison more complicated. The shifts are typically less than a pixel but can be more severe in some cases; our worst case is NGC3351, where the shift between NIRCam and MIRI is around one MIRI pixel. In a few cases there are multiple peaks in the distribution of residual astrometric offsets, indicating some persistent unresolved issues with the tile-to-tile matching.
@@ 5 Existing mosaic comparison
In this Section, we show the level of changes in our data processing relative to both our early science data release and the mosaics obtainable through the MAST service. Overall, our latest version (DR1.1.0) is a significant improvement over both of these products, and so we suggest users to prefer our latest versions over these other available versions.
@@ 5.1 Comparison to early science release
Figure 14: Comparison between our latest reprocessing and our early science mosaic for NGC 628. The difference here is the early science version subtracted from our latest reprocessed version, after reprojecting to the same WCS. There are significant differences in the short NIRCam band, caused by both background level matching and our new destriping method. At long NIRCam wavelengths, the differences are relatively minor. For MIRI, the improved calibration shows up as a galaxy imprint (most clear at F770W). Differences in astrometry show neighboring red/blue shifts and the background matching algorithm shows clear steps between mosaic tiles. The complete figure set (4 images) is available in the online journal.
@@
We will be revisiting many of our early science results with the full Cycle 1 sample in due course, as the increased number statistics and various changes to the data may affect results. For the most part, our First Results works took a conservative approach and many focused on source detection rather than detailed photometry. These results are unlikely to be affected by our improvements to the data processing. However, those that used flux measurements are likely to vary at the 10-20 percent level, especially those using long wavelength MIRI data, and any new publications will supersede the First Results (e.g. Sutter et al., subm). Initial testing has also shown that the improvements to astrometry allow for better measurements of the concentration index (CI, see Rodríguez et al., 2023). This means that in the F200W band we can better discriminate star clusters from individual stars (Rodríguez et al., in prep.).